OpenStudy (anonymous):

Can someone help me solve for x? 2x3+6x2-x = 3

OpenStudy (anonymous):

hint 2x3+6x2-x -3=0

OpenStudy (anonymous):

2x3+6x2-x -3=0 2x^2(x+3) -1(x+3)=0 or (x+3)(2x^2-1)=0 hence x=-3 or x=+- 1/sqrt(2)

OpenStudy (anonymous):

So, x could be either answers mat?

OpenStudy (anonymous):

yup